# https://yh2371.github.io/como

### v1

@@ 0.A.3 Fine-grained Keyword Setup
Fine-grained keywords
Head
The man’s head tilts slightly forward and turns to the right, following the direction of the stumble
Torso
The man’s torso leans to the right, as if losing balance.
L-Arm
The man’s left arm swings outward to the left, in an instinctive attempt to regain balance
R-Arm
The man’s right arm tucks in towards his body as he stumbles to the right
L-Hand
The man’s left hand is open, ready to grasp anything in the vicinity for support if needed
R-Hand
The man’s right hand clenches slightly, moving in unison with the right arm
L-Leg
The man’s left leg is firmly planted, acting as the pivot for the stumble
R-Leg
The man’s right leg lifts and steps awkwardly to the right, causing the stumble
L-Feet
The man’s left foot remains grounded, providing the only source of stability during the stumble
R-Feet
The man’s right foot lands unevenly on the ground, leading to the stumble
Mood
Unsteady
@@ 0.A.5 Motion Editing Prompts
Motion is represented by a set of joint states, defined as follows: Table 1 Joint State Meanings (Key: Joint State Index, Value: Joint State Meaning): {table1} Given the edit instruction: {edit} Return a semi-colon separated sequence of the ids of the joint states you will need to examine in order to determine the starting and ending frame of a motion sequence that will be affected by the edit instruction. Format example: 0;1;5;9. Do not reply anything else.
You will be provided with a text description of the motion, a motion code sequence and a motion edit instruction. You are be required to determine the starting and ending frame of the sequence that will be affected by the edit. Here is what you need to know about the encoding of the motion sequences: The motion is represented a number of time frames, each time frame contains a set of joint states, each joint state contains a code value. The definitions are: Table 1 Joint State Meanings (Key: Joint State Index, Value: Joint State Meaning): {table1} Table 2 Code Meaning (Key: Code ID, Value: Code Meaning): {table2} Rules: smaller angles indicates more bending. The motion code sequence is: {codes} The total number of time frames is {length} The text description is: {details} The edit instruction is: {edit} Return the starting index and ending index of the segment that is affected by the edit, separated by semi-colon, if the edit affects the overall movement, select the entire sequence. Format example: 0;19. Do not reply anything else.
@@
Motion is represented by a set of joint states, defined as follows: Table 1 Joint State Meanings (Key: Joint State Index, Value: Joint State Meaning): {table1} Given the edit instruction: {edit} Return a semi-colon separated sequence of the ids of the joint states you may be affected by the edit instruction. Format example: 0;1;5;9. Do not reply anything else.
You will be provided with a text description of the motion, a motion code sequence for a given joint state and a motion edit instruction. You will be required to determine how to modify the codes within the provided sequence accordingly. Here is what you need to know about the encoding of the motion sequences: The motion is represented as a list of joint states of length T, T is the number time frames. Each joint state contains a code value. The usable codes are defined as follows: Table 1 Usable Code Meaning (Key: Code ID, Value: Code Meaning): {table2} Rules: smaller angles indicates more bending. You are given this motion code sequence for the joint state {joint}, it has already been sliced to keep only the segment you will need to edit: {codes}. The text description of the overall motion sequence is: {details}. The edit instruction is: {edit} Return the edited motion only as a sequence of integer code ids of length {length} separated by semi-colons, only use code ids in the provided table. If no edit needs to be made, return the original sequence. Format example: 1;2;3;4. Do not reply anything else. No explanation needed.
@@ 0.A.6 Inference Time
Methods AITS (s) normal-$\downarrow$
@@ Appendix 0.B Human Evaluation
A person lowers their arms, and then moves them back up to shoulder height
Keep both knees deeply bent
A person, with both knees deeply bent, lowers their arms and then raises them back up to shoulder height
The person bent down and dodge something towards the left
Bend down slower
The person slowly bent down and dodged something towards the left
A person walks forward and then appears to bump into something, then continues walking forward
Make the bump more dramatic
A person walks forward, then suddenly collides with a large unseen obstacle with a significant impact, recoiling notably before resuming their forward motion
A person beginning to run in a straight line
Raise left hand at the end
A person begins to run in a straight line and raises their left hand near the end of the run
Table B10: Example updated descriptions for baseline motion editing.

### v2

@@ 5 Conclusion
Acknowledgment
This research is supported in part by the Office of the Director of National Intelligence (ODNI), Intelligence Advanced Research Projects Activity (IARPA), via the HIATUS Program contract #2022-22072200005. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of ODNI, IARPA, or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.
@@ 0.A.3 Fine-grained Keyword Setup
Fine-grained keywords Head The man’s head tilts slightly forward and turns to the right, following the direction of the stumble
Torso The man’s torso leans to the right, as if losing balance.
L-Arm The man’s left arm swings outward to the left, in an instinctive attempt to regain balance
R-Arm The man’s right arm tucks in towards his body as he stumbles to the right
L-Hand The man’s left hand is open, ready to grasp anything in the vicinity for support if needed
R-Hand The man’s right hand clenches slightly, moving in unison with the right arm
L-Leg The man’s left leg is firmly planted, acting as the pivot for the stumble
R-Leg The man’s right leg lifts and steps awkwardly to the right, causing the stumble
L-Feet The man’s left foot remains grounded, providing the only source of stability during the stumble
R-Feet The man’s right foot lands unevenly on the ground, leading to the stumble
Mood Unsteady
@@ 0.A.5 Motion Editing Prompts
Motion is represented by a set of joint states, defined as follows:
Table 1 Joint State Meanings (Key: Joint State Index, Value: Joint State Meaning): {table1}
Given the edit instruction: {edit}
Return a semi-colon separated sequence of the ids of the joint states you will need to examine in order to determine the starting and ending frame of a motion sequence that will be affected by the edit instruction.
Format example: 0;1;5;9. Do not reply anything else.
You will be provided with a text description of the motion, a motion code sequence and a motion edit instruction. You are be required to determine the starting and ending frame of the sequence that will be affected by the edit. Here is what you need to know about the encoding of the motion sequences:
The motion is represented a number of time frames, each time frame contains a set of joint states, each joint state contains a code value. The definitions are:
Table 1 Joint State Meanings (Key: Joint State Index, Value: Joint State Meaning): {table1}
Table 2 Code Meaning (Key: Code ID, Value: Code Meaning): {table2}
Rules: smaller angles indicates more bending.
The motion code sequence is: {codes}
The total number of time frames is {length}
The text description is: {details}
The edit instruction is: {edit}
Return the starting index and ending index of the segment that is affected by the edit, separated by semi-colon, if the edit affects the overall movement, select the entire sequence. Format example: 0;19. Do not reply anything else.
@@
Motion is represented by a set of joint states, defined as follows:
Table 1 Joint State Meanings (Key: Joint State Index, Value: Joint State Meaning): {table1}
Given the edit instruction: {edit}
Return a semi-colon separated sequence of the ids of the joint states you may be affected by the edit instruction. Format example: 0;1;5;9. Do not reply anything else.
You will be provided with a text description of the motion, a motion code sequence for a given joint state and a motion edit instruction. You will be required to determine how to modify the codes within the provided sequence accordingly.
Here is what you need to know about the encoding of the motion sequences: The motion is represented as a list of joint states of length T, T is the number time frames. Each joint state contains a code value. The usable codes are defined as follows:
Table 1 Usable Code Meaning (Key: Code ID, Value: Code Meaning): {table2}
Rules: smaller angles indicates more bending.
You are given this motion code sequence for the joint state {joint}, it has already been sliced to keep only the segment you will need to edit: {codes}.
The text description of the overall motion sequence is: {details}.
The edit instruction is: {edit}
Return the edited motion only as a sequence of integer code ids of length {length} separated by semi-colons, only use code ids in the provided table. If no edit needs to be made, return the original sequence. Format example: 1;2;3;4. Do not reply anything else. No explanation needed.
@@ 0.A.6 Inference Time
Methods AITS (s) $\downarrow$
@@ Appendix 0.B Human Evaluation
Table B10: Example updated descriptions for baseline motion editing.
@@
A person lowers their arms, and then moves them back up to shoulder height Keep both knees deeply bent A person, with both knees deeply bent, lowers their arms and then raises them back up to shoulder height
The person bent down and dodge something towards the left Bend down slower The person slowly bent down and dodged something towards the left
A person walks forward and then appears to bump into something, then continues walking forward Make the bump more dramatic A person walks forward, then suddenly collides with a large unseen obstacle with a significant impact, recoiling notably before resuming their forward motion
A person beginning to run in a straight line Raise left hand at the end A person begins to run in a straight line and raises their left hand near the end of the run